OpenStudy (anonymous):

A travel mug is shaped like a cylinder. It is 9 centimeters long and 15 centimeters tall. Find its volume to the nearest tenth. Use 3.14 for pi

OpenStudy (littlebird):

V=pir^2h

OpenStudy (littlebird):

r=9/2 so r^2=81/4 (I'm assuming by "long" you mean it has that diameter) h=15 pi=3.14

OpenStudy (anonymous):

so you mean it would be 3.14*\[3.14*4.5\prime/2*15?\]

OpenStudy (anonymous):

@Littlebird

OpenStudy (littlebird):

pi+(4.5)^2+15=38.39 This only works if I was correct in assuming that 9 cm is the diameter. The r stands for radius. Radius is half the diameter. That is why I was dividing 9 by 2 before I square it.